# FENDA-FL: Personalized Federated Learning on Heterogeneous Clinical Datasets

### v2

@@ Appendix A FLamby Details and Hyper-parameters
In this section, details are provided around each of the selected datasets in the FLamby benchmark. For more information on items such as preprocessing and feature engineering, see (du Terrail et al., 2022a). The library provided with the FLamby benchmark is used to replicate the training parameters, loss functions, model architectures, and other experimental components for consistency. Table LABEL:flamby_dataset_hp_studies details the hyper-parameter sweeps and best values used for the results reported in Section 4, which differ from those of the original FLamby settings. Note that the searched values for MOON and PerFCL are informed by the original papers. The best hyper-parameters are chosen by computing the lowest aggregated validation loss seen by the server during FL training, averaged over five training runs. The aggregated validation loss is computed by averaging the loss on each client’s validation set, weighted by the number of training examples held by each client. For central and local training, a setup identical to the FLamby benchmark is used. As a result, the best performing parameters optimized in the benchmark are applied for each task. Client dataset distributions for each task are shown in Table LABEL:fl_dataset_distributions.
In selecting the FL algorithms for the FLamby and GEMINI experiments, FedAdam is chosen as representative of the FedOpt family (Reddi et al., 2021) because it performed well in the original FLamby experiments. Similarly, the Cyclic method is not considered here as it performed poorly across all original FLamby benchmarks.
A.1 Fed-Heart-Disease
@@ A.2 Fed-IXI
A.3 Fed-ISIC2019
The Fed-ISIC2019 dataset is split across six different clients. Three of the client datasets are derived from the same source, the Medical University of Vienna, but were generated by different imaging devices. A weighted focal loss function (Lin et al., 2020) is used to fine-tune the EfficientNet-B0 architectures and balanced accuracy is measured to assess performance. All training runs use a batch size of $64$. Central and local training uses an AdamW optimizer with a learning rate of $5\mathrm{e}{\text{-}4}$ over $20$ epochs. As seen in Table LABEL:fl_dataset_distributions, the number of data-points held by each client varies dramatically.
[…]
gemini_dataset_hp_studies Task     Mortality Delirium Method Params. Range Best Range Best FedAvg Client LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$ FedAdam Client LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}4}$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}4}$ Server LR $1.0$ to $1\mathrm{e}{\text{-}3}$ $1\mathrm{e}{\text{-}3}$ $1.0$ to $1\mathrm{e}{\text{-}3}$ $1\mathrm{e}{\text{-}3}$ FedProx Client LR $1.0$ to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.1$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$ $\mu$ $1.0$ to $0.01$ $0.1$ $1.0$ to $0.01$ $0.1$ SCAFFOLD Client LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$ Server LR $1.0$ to $1\mathrm{e}{\text{-}3}$ $1.0$ $1.0$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$ MOON Client LR $0.1$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$ $0.1$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}3}$ $\mu$ $10$ to $1\mathrm{e}{\text{-}3}$, $5$ $10$ $10$ to $1\mathrm{e}{\text{-}3}$, $5$ $10$ FedPer Client LR $0.1$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$ $0.1$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}5}$ APFL Client LR $0.1$ to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.01$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$ $\alpha$ LR $1.0$ to $1\mathrm{e}{\text{-}3}$ $0.01$ $1.0$ to $1\mathrm{e}{\text{-}3}$ $0.01$ PerFCL Client LR $0.1$ to $1\mathrm{e}{\text{-}3}$ $1\mathrm{e}{\text{-}3}$ $0.1$ to $1\mathrm{e}{\text{-}3}$ $1\mathrm{e}{\text{-}3}$ $\mu$ $1.0$ to $0.01$ $0.1$ $1.0$ to $0.01$ $1.0$ $\gamma$$\{1,5,10\}$ $10$$\{1,5,10\}$ $1.0$ FENDA-FL Client LR $0.01$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$ $0.01$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$ Task     Mortality Delirium Method Params. Range Best Range Best Local 0 LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$ Local 1 LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$ Local 2 LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}4}$ Local 3 LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$ Local 4 LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$ Local 5 LR 0.01 to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$ Local 6 LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$- -- Central LR $0.01$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$ $0.01$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$
@@ A.3 Fed-ISIC2019
[…]
fl_dataset_distributions Dataset Number Train Size Test Size 0 199 104 Fed-Heart- 1 172 89 Disease 2 30 16 3 85 45 Fed-IXI 0 262 66 1 142 36 2 54 14 Fed-ISIC2019 0 9930 2483 1 3163 791 2 2691 672 3 1807 452 4 655 164 5 351 88 Dataset Number Train Size Test Size GEMINI Mortality 0 13234 3308 1 21018 5254 2 14180 3545 3 15251 3813 4 14441 3610 5 12718 3180 6 23820 5955 GEMINI Delirium 0 660 101 1 590 99 2 1218 199 3 494 76 4 574 109 5 1392 258
Table 7: Train and test set sizes across clients in the FLamby datasets, reproduced from (du Terrail et al., 2022a), and the GEMINI datasets. Each GEMINI client represents a separate Canadian institution.
Appendix B GEMINI Details and Hyper-parameters
@@
GEMINI is a research network in which hospitals share data with a central repository. As the repository contains detailed patient information, access to GEMINI data is granted only after approval, in accordance with the network’s data sharing agreement and research ethics protocols. Similar to the FLamby experiments, the best hyper-parameters are chosen based on the minimum average validation loss computed by the server. The loss is the weighted average of clients’ model’s loss in each round of FL training. The sweep space and best hyper-parameters found for each of the FL methods are reported in Table LABEL:gemini_dataset_hp_studies. Hyper-parameter sweeps are also performed for central and local model training for both GEMINI tasks, the results of which are also found in the table.
While both the Mortality and Delirium prediction tasks are drawn from GEMINI consortium data, the data pre-processing steps result in two distinct task-specific datasets. Table LABEL:fl_dataset_distributions displays the client data distribution for both tasks. An overview of the data preparation steps is provided in the following sections. The code may be shared with researchers upon receiving access to GEMINI.
B.1 Mortality Prediction
@@
For this task, there are seven participating hospitals, each representing a client with its own private data. A binary cross-entropy loss is used for training. For all experiments, the batch size is set to $64$. The number of samples for mortality prediction is large compared to other tasks. A careful analysis resulted in the selection of $35$ features that were subsequently processed by a standard data pipeline, which included data vectorization and scaling. A non-exhaustive list of selected features includes: readmission status, whether the patient is transferred from an acute care institution, length of stay in days, CD-10-CA classification code, counts of previous encounters, diagnosis type, and age. Categorical features require special preprocessing in a federated setting. For clients to consistently map such features to a numerical space, we assume that the server has oracle knowledge of all possible feature values and provides one-hot mappings to all clients.
B.2 Delirium Prediction
@@
For this task, data is drawn from six hospitals. Each hospital participates in FL as a single client. As seen in Table LABEL:fl_dataset_distributions, the number of samples in this task is smaller than in mortality prediction. This is because delirium labels are provided through a rigorous and time-intensive process of medical record review by a trained expert (Inouye et al., 2005). This is one of the only validated methods for such labeling, as it is poorly captured in routine administrative or clinical data. A wide variety of medical features is used, such as count of room transfer and ICU transfer, whether the patient has entered an intensive care unit since their admission, length of stay in days, laboratory-based Acute Physiology Score, ICU length of stay in hours, duration of the patient’s stay in the ER in hours, whether admission was via ambulance, triage level, total number of transfusions, type of blood product or component transfused, medication names, text description of radiology images, among many others. The feature engineering pipeline results in $1,119$ features.
@@
Concretely, the data pre-processing steps are the following. One-hot encoding is used for categorical features such as gender, discharge disposition, and admission category. A TF-IDF vectorizer is used to vectorize text-based columns such as radiology image and medication descriptions. A MaxAbsScaler is used for data scaling. Finally, during training, resampling is performed to balance negative and positive samples and increase the dataset size. Resampling is done separately in each client based on the client’s data. Note that resampling is not performed on the test datasets. Similar to the mortality task, we also assume oracle knowledge in the server, which enables clients to map categorical features into the same space. If clients individually process such features, there is a high chance of misalignment, resulting in heterogeneous feature spaces. Finally, during training, the batch size is $64$ for all methods.
Appendix C Trainable Parameters
Table LABEL:experimental_model_parameters provides a comparison of the trainable parameters present in the models used in the FLamby and GEMINI experiments. The “standard” column refers to models trained in all settings, excluding APFL, FENDA-FL, and PerFCL. Note that both FedPer and MOON use the $151$ parameter models for Fed-Heart-Disease to facilitate feature extraction. For a fair comparison, we have sought to keep the number of trainable parameters across different architectures close to ensure that models for different methods do not have out-sized expressiveness.
[…]
experimental_model_parameters Dataset Standard APFL FENDA-FL PerFCL Fed-Heart-Disease $14$ or $151$ $152$ $151$ $151$ Fed-IXI $1,106,520$ $984,624$ $984,620$ $984,620$ Fed-ISIC2019 $4,017,796$ $4,631,088$ $4,775,016$ $4,775,016$ Mortality $222,784$ $222,784$ $222,784$ $222,784$ Delirium $8,988,736$ $8,988,736$ $8,717,448$ $8,717,448$
Table 8: Trainable parameter counts for the model architectures used in the FLamby and GEMINI experiments. Standard refers to models used in all settings outside of APFL, FENDA-FL, and PerFCL.
[…]
fed_heart_disease_all_models Refer to caption
Figure 3: Mean accuracy for the Fed-Heart-Disease task. This figure shows results for small and large models with $14$ and $151$ parameters, respectively, for strategies other than FENDA-FL and APFL. Strategies sub-scripted with an S use small models and those with an L use the large models. Solid and hatched bars indicate global and local checkpointing, respectively.
Appendix D Fed-Heart-Disease: Large Models
Figure LABEL:fed_heart_disease_all_models displays the performance of both small and large models for various training strategies on Fed-Heart-Disease. The small models are logistic regression with only $14$ trainable parameters. The large versions are two-layer DNNs with $151$ parameters. The models used by FENDA-FL/PerFCL and APFL are DNN-based with $151$ and $152$ parameters, respectively. For strategies other than FENDA-FL, PerFCL, and APFL, the large model variants generally do not improve performance. Note that federated checkpointing, as described in Section 3.3, is applied in all cases. Therefore, it is unlikely that this phenomenon is due to over-fitting.
Appendix E FENDA-FL Architecture Studies
As discussed in Section 3.1, an important advantage of the FENDA-FL setup is that the local and global feature extraction modules need not be symmetric with equal latent space sizes. This allows for the injection of inductive bias through architecture choice. For example, in situations where global models perform well, one may assign more parameters and a larger latent space to FENDA-FL’s global feature extractor. In this section, we show that such choices can materially boost performance.
[…]
delirium_architecture_perturbations Layers (G, L) Latent (G, L) Parameters Acc. AUC/ROC $4$, $4$ $64$, $64$ $8,631,424$ $0.7926$ ($0.0071$) $0.8222$ ($0.0036$) $4$, $2$ $64$, $64$ $8,729,600$ $0.8024$ ($0.0037$) $0.8368$ ($0.0039$) $4$, $2$ $128$, $8$ $8,717,448$ $\mathbf{0.8064}$ ($0.0036$) $\mathbf{0.8518}$ ($0.0017$) $2$, $4$ $64$, $64$ $8,729,600$ $0.7935$ ($0.0073$) $0.8153$ ($0.0059$) $2$, $4$ $8$, $128$ $8,717,448$ $0.7935$ ($0.0073$) $0.8153$ ($0.0059$)
@@ Appendix E FENDA-FL Architecture Studies
Table 9: Results corresponding to different FENDA-FL architecture variations on the Delirium task. “Layers (G, L)” and “Latent (G, L)” refer to the number of dense layers assigned to the global and local feature extractors, respectively, along with the size of the corresponding latent spaces.

### v3

@@ Appendix D FLamby Details and Hyper-parameters
In this section details are provided for the datasets selected from the FLamby benchmark. For more information on items such as preprocessing and feature engineering, see (du Terrail et al., 2022a). The library provided with the benchmark is used to replicate the training parameters, loss functions, model architectures, and other experimental components for consistency. Table 7 details the hyper-parameter sweeps and best values used for the results reported in Section 5, which differ from the original FLamby settings. Note that the searched values for MOON, PerFCL, and Ditto are informed by the original papers. The best hyper-parameters are chosen by computing the lowest aggregated validation loss seen by the server during FL training, averaged over five training runs. The aggregated loss is computed by averaging each client’s validation loss, weighted by the number of examples held by each client. For central and local training, a setup identical to the FLamby benchmark is used. As a result, the best performing parameters optimized in the benchmark are applied for each task. Client dataset sizes for each task are shown in Table 8.
Table 8: Train and test set sizes across clients in the FLamby datasets, reproduced from (du Terrail et al., 2022a), and GEMINI datasets. Each GEMINI client represents a separate Canadian institution.
Dataset Number Train Size Test Size
0 199 104
Fed-Heart- 1 172 89
Disease 2 30 16
3 85 45
Fed-IXI 0 262 66
1 142 36
2 54 14
Fed-ISIC2019 0 9930 2483
1 3163 791
2 2691 672
3 1807 452
4 655 164
5 351 88
Dataset Number Train Size Test Size
GEMINI Mortality 0 13234 3308
1 21018 5254
2 14180 3545
3 15251 3813
4 14441 3610
5 12718 3180
6 23820 5955
GEMINI Delirium 0 660 101
1 590 99
2 1218 199
3 494 76
4 574 109
5 1392 258
Fed-Heart-Disease:
@@
Fed-ISIC2019:
The Fed-ISIC2019 dataset is split across six clients. Three of the client datasets are derived from the same source, the Medical University of Vienna, but were generated by different imaging devices. A weighted focal loss function (Lin et al., 2020) is used to fine-tune EfficientNet-B0 architectures and balanced accuracy is measured to assess performance. All training runs use a batch size of $64$. As seen in Table 8, the number of data points held by each client varies dramatically.
D.1 Sub-Optimal Hyper-Parameters For Checkpointing Study
For all strategies, except PerFCL, a client-side learning rate of $0.1$ is applied. For FedAdam and SCAFFOLD, a server-side learning rate of $0.1$ and $1.0$, respectively, is used. FedProx sets $\mu$ = $1\mathrm{e}{\text{-}3}$, APFL applies a learning rate for $\alpha$ of 1.0, MOON sets $\mu$ to $1\mathrm{e}{\text{-}3}$, and Ditto uses $\lambda=0.01$. Finally, PerFCL uses a client-side learning rate of $1\mathrm{e}{\text{-}3}$, $\mu=0.01$, and $\gamma=1$.
@@ D.1 Sub-Optimal Hyper-Parameters For Checkpointing Study
Task     Mortality Delirium
Method Params. Range Best Range Best
FedAvg Client LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$
FedAdam Client LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}4}$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}4}$
Server LR $1.0$ to $1\mathrm{e}{\text{-}3}$ $1\mathrm{e}{\text{-}3}$ $1.0$ to $1\mathrm{e}{\text{-}3}$ $1\mathrm{e}{\text{-}3}$
FedProx Client LR $1.0$ to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.1$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$
$\mu$ $1.0$ to $0.01$ $0.1$ $1.0$ to $0.01$ $0.1$
SCAFFOLD Client LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$
Server LR $1.0$ to $1\mathrm{e}{\text{-}3}$ $1.0$ $1.0$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$
MOON Client LR $0.1$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$ $0.1$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$
$\mu$ $10$ to $1\mathrm{e}{\text{-}3}$, $5$ $1\mathrm{e}{\text{-}3}$ $10$ to $1\mathrm{e}{\text{-}3}$, $5$ $1.0$
FedPer Client LR $0.1$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$ $0.1$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}5}$
Ditto Client LR $0.1$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$ $0.1$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}5}$
$\lambda$ $1.0$ to $0.01$ $0.01$ $1.0$ to $0.01$ $0.01$
APFL Client LR $0.1$ to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.01$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$
$\alpha$ LR $1.0$ to $1\mathrm{e}{\text{-}3}$ $0.01$ $1.0$ to $1\mathrm{e}{\text{-}3}$ $0.01$
PerFCL Client LR $0.1$ to $1\mathrm{e}{\text{-}3}$ $1\mathrm{e}{\text{-}3}$ $0.1$ to $1\mathrm{e}{\text{-}3}$ $1\mathrm{e}{\text{-}3}$
$\mu$ $1.0$ to $0.01$ $0.1$ $1.0$ to $0.01$ $1.0$
$\gamma$$\{1,5,10\}$ $10$$\{1,5,10\}$ $1.0$
FENDA-FL Client LR $0.01$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$ $0.01$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$
Task     Mortality Delirium
Method Params. Range Best Range Best
Local 0 LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$
Local 1 LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$
Local 2 LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}4}$
Local 3 LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$
Local 4 LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$
Local 5 LR 0.01 to $1\mathrm{e}{\text{-}4}$ $0.01$ $0.01$ to $1\mathrm{e}{\text{-}4}$ $1\mathrm{e}{\text{-}3}$
Local 6 LR $0.01$ to $1\mathrm{e}{\text{-}4}$ $0.01$- --
Central LR $0.01$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$ $0.01$ to $1\mathrm{e}{\text{-}5}$ $1\mathrm{e}{\text{-}4}$
Appendix E GEMINI Details and Hyper-parameters
@@
GEMINI is a research network in which hospitals share data with a central repository. As the repository contains detailed patient information, access to GEMINI data is granted only after approval, in accordance with the network’s data sharing agreement and research ethics protocols. As in the FLamby experiments, hyper-parameters are chosen based on the minimum weighted-average validation loss computed by the server. The sweep space and best hyper-parameters for each method are reported in Table 9. Hyper-parameter sweeps are also performed for central and local training for the GEMINI tasks, the results of which are also found in the table. In addition, Table 8 displays the client data distributions. An overview of the data preparation steps is provided in subsequent sections. The code is available to be shared with researchers upon receiving access to GEMINI.
Federated Pre-Processing
Data alignment in FL pipelines is challenging. For example, categorical features require special attention in a federated setting. For the GEMINI tasks, a federated pre-processing step is included for such features to synchronize vectorization and ensure that the resulting feature spaces remain aligned across clients. For clients to consistently map such features to a numerical space, we assume that the server has oracle knowledge of all possible feature values and provides one-hot mappings to all clients.
E.1 Mortality Prediction
@@
For this task, there are seven hospitals, each representing a client with its own private data. A binary cross-entropy loss is used for training, with a batch size of $64$. The number of samples for mortality prediction is large compared to other tasks. A careful analysis resulted in the selection of $35$ features that were subsequently processed by a standard data pipeline, which included data vectorization and scaling. For more details about the data pre-processing, see (Subasri et al., 2023). A non-exhaustive list of features includes: readmission status, whether the patient is transferred from an acute care institution, length of stay (LOS) in days, CD-10-CA classification code, counts of previous encounters, diagnosis type, and age. Patient characteristics are described in Table 10 for a subset of features.
Table 10: Feature characteristics for the Mortality prediction dataset conditioned on label.
Mortality Label False
Client 0 1 2 3 4 5 6
Patients 15122 22719 16127 16645 16250 14083 27448
Avg. Prev. Encounters 0.68 0.59 1.19 0.72 0.82 0.89 0.59
Avg. LOS (Days) 8.71 10.08 7.89 8.46 7.54 9.02 8.43
% Emergent Triage 60.55 46.83 50.88 34.71 50.01 42.7 46.11
% Nursing Home 10.5 11.83 4.19 6.41 4.48 10.23 8.29
Mortality Label True
Client 0 1 2 3 4 5 6
Patients 1420 3553 1598 2419 1801 1815 2327
Avg. Prev. Encounters 1.36 1.16 1.53 1.28 1.31 1.62 1.11
Avg. LOS (Days) 18.96 17.66 17.58 14.52 12.74 14.06 17.03
% Emergent Triage 57.81 47.62 55.13 44.85 59.57 54.15 49.46
% Nursing Home 30.77 29.63 13.82 12.48 8.21 22.09 17.61
E.2 Delirium Prediction
@@
For this task, data is drawn from six hospitals, each participating in FL as a distinct client. As seen in Table 8, the number of samples in this task is smaller than in mortality prediction. This is because delirium labels are provided through a rigorous and time-intensive process of medical record review by a trained expert (Inouye et al., 2005). This is one of the only validated methods for such labeling, as it is poorly captured in routine administrative or clinical data. A wide variety of medical features is used, such as count of room transfer and ICU transfer, whether the patient has entered an intensive care unit since their admission, length of stay in days, laboratory-based Acute Physiology Score, ICU length of stay in hours, duration of the patient’s stay in the ER in hours, whether admission was via ambulance, triage level, total number of transfusions, type of blood product or component transfused, medication names, text description of radiology images, among many others. The feature engineering pipeline results in $1,119$ features.
@@
The data pre-processing steps are the following. One-hot encoding is used for categorical features such as gender, discharge disposition, and admission category. A TF-IDF vectorizer is used to vectorize text-based columns such as radiology image and medication descriptions. A MaxAbsScaler is used for data scaling. Finally, during training, resampling is performed to balance negative and positive samples and increase the dataset size. Resampling is done separately in each client based on the client’s data. Note that resampling is not performed on the test datasets. Similar to the mortality task, we assume oracle knowledge in the server, enabling clients to map categorical features into a shared space. A batch size of $64$ is used for all methods. In contrast to the Mortality dataset, the patient population for Delirium is fairly small. Therefore, no further statistics or patient characteristics are provided to prevent possible leakage of private information.
@@
Table 11: Results corresponding to different FENDA-FL architecture variations on the Delirium task. “Layers (G, L)” and “Latent (G, L)” refer to the number of dense layers assigned to the global (G) and local (L) feature extractors, respectively, along with the size of the corresponding latent spaces.
Layers (G, L) Latent (G, L) Parameters Acc. AUC/ROC
$4$, $4$ $64$, $64$ $8,631,424$ $0.7926$ ($0.0071$) $0.8222$ ($0.0036$)
$4$, $2$ $64$, $64$ $8,729,600$ $0.8024$ ($0.0037$) $0.8368$ ($0.0039$)
$4$, $2$ $128$, $8$ $8,717,448$ $\mathbf{0.8064}$ ($0.0036$) $\mathbf{0.8518}$ ($0.0017$)
$2$, $4$ $64$, $64$ $8,729,600$ $0.7935$ ($0.0073$) $0.8153$ ($0.0059$)
$2$, $4$ $8$, $128$ $8,717,448$ $0.7935$ ($0.0073$) $0.8153$ ($0.0059$)
Appendix F FENDA-FL Architecture Studies
As discussed in Section 3.3, an important advantage of FENDA-FL is that the local and global feature extraction modules need not be symmetric with equal latent space sizes. This allows for the injection of inductive bias through architecture choice. For example, in situations where global models perform well, one may assign more parameters and a larger latent space to FENDA-FL’s global feature extractor. In this section, we show that such choices can materially boost performance.